OpenStudy (anonymous):

antiderivative of 2sec^2x-sinx+2e^x

OpenStudy (anonymous):

These integrals should be things that you've memorized by now; the derivative of tan(x) = sec^2(x). The others shouldn't be difficult at all; just break them up into 3 separate ones.

OpenStudy (anonymous):

Alright so we have:\[2\int\limits_{}^{}\sec^2 xdx - \int\limits_{}^{}sinxdx + 2\int\limits_{}^{}e^xdx\]\[= 2tanx + cosx + 2e^x + c\] ^_^

OpenStudy (anonymous):

is it clear now? :)

OpenStudy (anonymous):

yes thanks! we actually haven't covered integrals yet. i just wanted to make sure i'm doing them right